OpenStudy (anonymous):

Blackfoot High School graduated 150 students. Approximately 15% of the senior class did not graduate. How many seniors were there total? I said 173 because 150*.15= 22.5 so then 150+22.5= 72.5 so approximately it would be 173 right?

OpenStudy (anonymous):

15% no graduate, means 85% graduated 85% = 150 1% = 150/85 100% = (150/85) x 100 = 176.47, but you cannot have 0.47 persons, so roudned down to 176

OpenStudy (anonymous):

100% = 100 units = total number of students 15% = 15 units = number of students that did not graduate 100%-15% = 85% = 85 units = number of students that graduated If 150 students graduated, this represents 85 units. Find 100 units to find total number of students 85 units = 150 1 unit = 150 / 85 = 1.764705882 100 units = 176.4705

jimthompson5910 (jim_thompson5910):

let x = total number of seniors 15% did not graduate, so 100 - 15 = 85% did graduate 85% of x = 0.85x = number who graduated so this means 0.85x = 150 x = 150/0.85 x = 176.470588235294 round this to the nearest whole number to get 176 so there are roughly 176 seniors